Can you solve for the following? Thanks!


[A financial analyst has been follow ng Fast Start Inc a new high-growth company. She estim ates that the current nsk-free rate is 6.25%, the market risk premium is 5% Fast Starts beta is 1.75. The current earnings per share (EPS ) is $2.50. The company has a 40% payout ratio. The analyst estimates that the companys dividend wil at a rate of 5% this year, 20% next year, and i 5% the following year After three years the dividend is expected to grow at a constant rate of 7% a year. The company to maintain its current payout ratio. The analyst believes that the stock is fairly priced. What is the current price of the stock?
Answer #1

Cost of equity = Risk free rate + Beta* Risk premium

k= 6.25%+ 1.75*5%

k=15%

D0 = EPS0* Dividend payout

= 2.5* 40% = $1

D1 = D0*(1+g) = 1*125% = 1.25

D2 = D1*(1+g) = 1.25*120%= 1.5

D3 = D2*(1+g) = 1.5*115% = 1.725

D4 = D3*(1+g)= 1.725*107% = 1.85

Horizon value = P3 = D4/(k-g)

=1.85/(15%-7%)

= 23.07

Current price = D1/(1+k)+ D2/(1+k)^2 + (D3+P3) /(1+k)^3

= 1.25/1.15+ 1.5/1.15^2 + (1.725+23.07) /1.15^3

=$18.52